OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y?

OpenStudy (merchandize):

2*4-3y=-10 8-3y=-10 -3y=-10-8 -3y=-18 y=6

OpenStudy (anonymous):

thanks

OpenStudy (merchandize):

pleasure